OpenStudy (anonymous):

can someone help me with this? find the first 5 derivatives of the function f(x)=e^x cosx, then find the 41st derivative. If possible, please show the work, thanks :)

OpenStudy (precal):

Use the product rule. Looks like your teacher wanted to keep you busy.....

OpenStudy (jamesj):

Cute question. What are the first four derivatives? We can detect a pattern; the idea obviously is not to differentiate 41 times.

OpenStudy (jamesj):

What's f'(x)?

OpenStudy (precal):

You still have to use the product rule..

OpenStudy (anonymous):

But is there a short cut to get to the 41st derivative? I mean, my professor can't seriously expect me to do this :(

OpenStudy (precal):

Do the first four .........

OpenStudy (jamesj):

Yes there is. That's why s/he asked you to do the first 4 first. So what are they?

OpenStudy (anonymous):

To be honest, I'm not even completely sure how to do those either. I have an exam tmrw, so I'm trying to understand as much as I can...

OpenStudy (jamesj):

If f(x) = e^x sin x, what is f'(x), just the first derivative?

OpenStudy (jamesj):

product rule

OpenStudy (jamesj):

sorry, let's stay with your function, f(x) = e^x . cos x

OpenStudy (anonymous):

ok,so derivative of cosx is -sinx, right?

OpenStudy (jamesj):

Yep, and I hope you've got that engraved in your brain for tomorrow!

OpenStudy (anonymous):

Uh, actually no. Sad to say I had to look that up, lol...

OpenStudy (precal):

First derivative is \[e ^{x}cosx-sinx(e ^{x})\] Factor out \[e ^{x}\] \[e ^{x}(cosx-sinx)\] You need to do your homework, because you can't wing calculus..

OpenStudy (anonymous):

I know, I don't know how I even got into this class. Math is my worst subject :(

OpenStudy (precal):

sorry, talk to your teacher...... :( I wish I could tell you more....

OpenStudy (anonymous):

Yea, I wish I could, but he just makes everything sound more complicated. He gives us the question and the answer, but nothing in between, so idk how to do the work.

OpenStudy (jamesj):

f'(x) = e^x (cos x - sin x) = f(x) - e^x sin x So f''(x) = f'(x) - e^x (sin x + cos x) = e^x ( cos x - sin x - sin x - cos x) = - 2 e^x sin x Now f'''(x) = - 2 e^x (sin x + cos x) f''''(x) = -2 e^x (sin x + cos x + cos x - sin x) = -4 e^x cos x = -4 f(x)

OpenStudy (jamesj):

Thus \[f^{(40)}(x) = (-4)^{10} f(x) = 2^{20} f(x)\] and \[f^{(41)}(x) = 2^{20} f'(x) = 2^{20}e^x(\cos x - \sin x)\]

OpenStudy (anonymous):

but i don't understand how for the 4th derivative, why did the -2 become -4?

OpenStudy (anonymous):

thats weird, looking back at the answer my professor gave me, he wrote that the 41st derivative is (-4^10)(-e^sinx+e^xcosx)

OpenStudy (jamesj):

What I have written is exactly equal.

OpenStudy (anonymous):

Oh, really? Well I really liked how you showed your steps. Are you a teacher or something?

OpenStudy (jamesj):

I was

OpenStudy (anonymous):

where did you teach?